OpenStudy (anonymous):

pleease help!! What is the solution set of the following equation? 1/5x^2 = x - 4/5 {-1, -4} {1, -4} {1, 4}

OpenStudy (mathmate):

You can substitute each solution set and see which one works. It is faster than solving a cubic.

OpenStudy (mathmate):

Don't know if you have a transcription error, the equation should read: (1/5)x^2 = x - 4/5

OpenStudy (anonymous):

so i substitute by putting the solution sets into the x place?

OpenStudy (anonymous):

i tried that, but maybe i'm doing it wrong, because it doesnt seem to make sence

OpenStudy (mathmate):

The equation should read : (1/5)x^2 = x - 4/5 to make sense.

OpenStudy (anonymous):

yeah, the equation already does read that, how do i substitute the solution sets?

OpenStudy (mathmate):

For example, if the solution set is {1,-4} put x=1 and see if the equation balances. If so, try x=-4. If both work, then you've found the solution set. Example, x=1: (1/5)(1)^2 =(1) - 4/5 See if the equation is valid, and so on.

OpenStudy (anonymous):

i keep getting .04=3.2

OpenStudy (zzr0ck3r):

how is this cubic?

OpenStudy (zzr0ck3r):

(1/5)x^2-x+4/5 = 0 x^2-5x+4 = 0 x^2 - 1x - 4x + 4 = 0 x(x-1)-4(x-1) = 0 (x-1)(x-4) = 0 x = 1 or x = 4